OpenStudy (anonymous):

can you factor out 2t^2-9t-12 ?

jimthompson5910 (jim_thompson5910):

no you cannot since there are no two numbers that multiply to 2(-12) = -24 AND that add to -9

OpenStudy (anonymous):

use the quadratic formula.

OpenStudy (anonymous):

okay i thought so...what math are you in ?

OpenStudy (anonymous):

Calc 1

OpenStudy (anonymous):

have you learned higher derivatives and the applications of the derivatives to velocity, acceleration, and rates of change ?

OpenStudy (anonymous):

Yeah i have

OpenStudy (anonymous):

Im not good at change of rate problems though :(

OpenStudy (anonymous):

rate of change**

OpenStudy (anonymous):

darn -.-

OpenStudy (anonymous):

I hate those with a passion lol.. i know how to find velocity and acceleration though.

jimthompson5910 (jim_thompson5910):

often, drawing a pic (a good one) really helps

OpenStudy (anonymous):

would you like to help me ?..possibly ?

jimthompson5910 (jim_thompson5910):

since you can lay out what the problem is giving you and what you want

jimthompson5910 (jim_thompson5910):

sure

OpenStudy (anonymous):

Velocity is the first derivative and acceleration is the second derivative

OpenStudy (anonymous):

the position function of a particle is given by s=t^3-4.5t^2-7t where t is greater than or equal to 0 a. when does the particle reach a velocity of 5 m/s b. when is the velocity 0 c. when is the acceleration 0

jimthompson5910 (jim_thompson5910):

the velocity function is simply the derivative of the position function

jimthompson5910 (jim_thompson5910):

so to answer parts a) and b), we need the velocity function

jimthompson5910 (jim_thompson5910):

so what do you get when you derive the position function s(t)?

OpenStudy (anonymous):

2t^2-9t-7

jimthompson5910 (jim_thompson5910):

good

jimthompson5910 (jim_thompson5910):

part a) when does the particle reach a velocity of 5 m/s This is just asking for what values of t make the velocity function equal to 5 Mathematically, they want to know which t values satisfy the equation v(t) = 5 But you found that v(t) = 2t^2-9t-7 So you're basically solving 2t^2-9t-7 = 5 for t

jimthompson5910 (jim_thompson5910):

2t^2-9t-7 = 5 2t^2-9t-7 - 5 = 0 2t^2-9t-12=0 You can't factor this, but we can solve using the quadratic formula t = (9+-sqrt(9^2-4(2)(-12)))/(2*2) t = (9+-sqrt(177))/(4) t = (9+sqrt(177))/(4) or t = (9-sqrt(177))/(4) t = 5.576 or t = -1.076 Throw out the negative solution So the only solution is t = 5.576 (this is a rough approximation of course) So the velocity is 5 m/s at approximately 5.576 seconds

jimthompson5910 (jim_thompson5910):

do the same for part b), but instead of using 5, use 0

OpenStudy (anonymous):

thank youu (:

jimthompson5910 (jim_thompson5910):

yw

OpenStudy (anonymous):

And then for acceleration you find the derivative of the velocity equation and set it equal to 0 and solve.

OpenStudy (anonymous):

okayy(: thank youu !